# Randomly Sampled Language Reasoning Problems Reveal Limits of LLMs

### v3

@@ Abstract
Can LLMs pick up language structure from examples? Evidence in prior work seems to indicate yes, as pretrained models repeatedly demonstrate the ability to adapt to new language structures and vocabularies. However, this line of research typically considers languages that are present within common pretraining datasets, or otherwise share notable similarities with seen languages. In contrast, in this work we attempt to measure models’ language understanding capacity while circumventing the risk of dataset recall. We parameterize large families of language tasks recognized by deterministic finite automata (DFAs), and can thus sample novel language reasoning problems to fairly evaluate LLMs regardless of training data. We find that, even in the strikingly simple setting of 3-state DFAs, LLMs underperform unparameterized $n$-Gram models on both language recognition and synthesis tasks. These results suggest that LLMs struggle to match the ability of basic language models in recognizing and reasoning over languages that are sufficiently distinct from the ones seen at training time, underscoring the distinction between learning individual languages and possessing a general theory of language.
@@ 1 Introduction
Figure 1: We use (a) randomly generated deterministic finite automata (DFAs) to sample previously unseen grammars to test models with. The DFA shown here, modeling the sum modulo 3 operation (with abc representing 0, 1, and 2 respectively), can be used to accept or reject strings from a 3-character alphabet. Accepted strings belong to the grammar, and rejected strings do not. We evaluate models on their ability to (b) act as a transducer, recognizing strings that belong to the DFA-defined grammar, and (c) generate new strings following the grammar.
@@
To circumvent this problem, we propose the following general approach: first define a large, exhaustive, and parsimoniously-defined space of languages that represents all languages of a certain difficulty level. Then, sample random languages from this space. By sampling randomly, we can guarantee no bias towards canonical languages that might share structure with ones in the training dataset. In this work, we use languages recognized by 3-state DFAs as these are the lowest nontrivial difficulty level, but this technique can be generalized to produce benchmarks of any difficulty level.
@@ 3.1 DFAs and Regular Languages
The original Chomsky Hierarchy (Chomsky, 1959) separates language into four types (Figure 2). We focus on the task of understanding Type 3 languages, the simplest form of language in the hierarchy, that are recognized by a Deterministic Finite Automaton (DFA). Examples of languages recognized by DFAs include simple ones like binary strings with an even number of ones, and even such examples as numbers in base 10 divisible by 7. Finite automaton languages are equivalent to regular languages, which are recognized by regular expressions.
One simple metric of the difficulty of a regular language is the number of states in the corresponding DFA, which represents the amount of memory the automaton has at any point while processing a given sequence. [^1] There are other metrics of difficulty, but we choose number of states as it is highly parsimonious. DFAs with 2 states have the property that their set of states is no larger than the output set $\{0,1\}$, and, therefore, do not have any hidden state. We thus explore 3-state DFAs, as this is the simplest nontrivial case.
@@ 3.2 Sequence Completion Task
We first pose a sequence completion task, in which models must complete a sequence in a given DFA’s language. We study this task primarily because, in practice, most language data that models encounter will be in roughly this format, with several example sequences in a given language followed by a distinct prefix that needs to be completed via next token prediction.
@@
We evaluate models by (1) sampling a DFA, (2) sampling 30 problem instances at random (each of which contains 30 example sequences and a distinct prefix), and then (3) computing a binary prediction score (whether or not the predicted completion creates a valid string in the language) for each instance separately. We then average this metric over several sampled DFAs to produce our accuracy score.
@@ 3.3 Transducer Task
In this task, an input sequence is annotated with an output at each token, the final output is masked, and the masked output is predicted by a language model. We call this a transducer task, as the DFA is used as a machine that converts a sequence of inputs into a sequence of outputs. E.g., given the language does the string have an even number of ’a’ tokens and the input abcabcaabbccaa, the annotated string (all that is provided to the model) is a0b0c0a1b1c1a0a1b1b1c1c1a0a and the output to predict is 1. For each problem instance, we provide 30 symbols, and for the first 29, the corresponding transducer output.
This task is significantly more transparent than the sequence completion task as the model has access to intermediate transducer outputs, an (imperfect) proxy for intermediate state. n
@@ 4 Experiments
We produce versions of each of these prompts for each task, denoting these with a subscript S for sequence completion prompts and T for transducer prompts. Full listings of these prompts can be found in Appendix E, with Table 5 containing a summary of each prompt. While no finite set of prompts will be fully sufficient to capture all possible model behavior, we believe our set of prompts captures common prompting strategies.
@@ 5.1 Sequence Completion
One potential problem with using this task for cross-model comparisons is the relevance of tokenization. Unfortunately, we found that forcing uniform tokenization by using commas in the prompt uniformly reduced accuracy, see Appendix D for details.
@@ 5.2 Transducer
Unlike sequence completion, this task is generally not fully determined, with the $\textsc{BruteForce}_{T}$ model only achieving 96.4% accuracy. However, relative results should still be valid as the impossible problem instances are present with equal probability for all models.
@@
Overall, we conclude that LLMs are unable to perform the DFA transducer inference task to a reasonable degree. This failure cannot be attributed to a lack of world modeling ability, as $n$-GramT models do not construct world models. Instead, it seems the LLMs are unable to detect patterns when those patterns are drawn from an unfamiliar source, even a relatively simple one.
@@ 5.4 Case Study: Sum Modulo 3 DFA
Total 30 30
a is no-op 21 22
1b and 1c lead to 0 14 17
2-periodic 9 14
3-periodic 4 4
2 red rooms 2 3
Figure 4: Results on Sum Modulo 3 DFA. (a) Venn diagram of errors (out of 1000). Labeled percentages are accuracies. (b) Results of qualitative analysis, out of 30 in both cases.
@@
Figure 4a depicts the number of errors each model receives on 1000 instances of the transducer task for this DFA. In general, nearly all errors made by the 6-GramT model were also made by at least one LLM model, while the two LLM models often made unique errors. Interestingly, while this task is better-known than most DFAs, we find that all 3 models perform worse on this DFA than their average across random DFAs.
We also performed a qualitative analysis, investigating CR’s outputs on the Red-GreenT prompt to see what kind of reasoning it is using; specifically we sampled 30 examples where it had the correct answer, and 30 examples where it had the incorrect answer but the 6-GramT model had the correct answer. Results of this analysis can be found in Figure 4b. We find that, in general, CR is following a 3-Gram approach, learning rules relating to the conditions under which the previous output and symbol can be used to predict the next output. Specifically, it is able to learn that a does not change the output, and that b and c will lead a 1 state to a 0 state. These results comport with the overall finding of Table 1, where we found that 3-GramT was the largest $n$-GramT that any LLM outperformed.
@@ Appendix B Results of o1-preview
Appendix C More details on Sum Modulo 3 DFA case study
@@
Appendix D Sequence Completion task prompt with Commas
@@
Appendix E Prompt Listings
@@ E.1 Summaries
E.2 Full example listings
E.2.1 BasicT

### v4

@@ Abstract
Can LLMs pick up language structure from examples? Evidence in prior work seems to indicate yes, as pretrained models repeatedly demonstrate the ability to adapt to new language structures. However, this line of research typically considers languages that are present within common pretraining datasets, or otherwise share notable similarities with seen languages. In contrast, in this work we attempt to measure models’ language understanding capacity while circumventing the risk of dataset recall. We parameterize large families of language tasks recognized by deterministic finite automata (DFAs), and can thus sample novel language reasoning problems to fairly evaluate LLMs regardless of training data. We find that, even in the strikingly simple setting of 3-state DFAs, LLMs underperform unparameterized $n$-Gram models on both language recognition and synthesis tasks. These results suggest that LLMs struggle to match the ability of basic language models in recognizing and reasoning over languages that are sufficiently distinct from the ones seen at training time, underscoring the distinction between learning individual languages and possessing a general theory of language.
@@ 1 Introduction
Figure 1: We sample randomly generated languages to test LLMs by sampling deterministic finite automata (DFAs). (a) The DFA shown here, modeling the sum modulo 3 operation (with abc representing 0, 1, and 2 respectively), can be used to accept or reject strings from a 3-character alphabet. Accepted strings belong to the grammar, and rejected strings do not. We evaluate models on their ability to (b) act as a transducer, recognizing strings that belong to the DFA-defined grammar, and (c) generate new strings following the grammar.
@@
To circumvent this problem, we propose the following general approach: first we define a large, exhaustive, and parsimoniously-defined space of languages that represents all languages of a certain difficulty level. Then, we sample random languages from this space. By sampling randomly, we can guarantee no bias towards canonical languages that might share structure with common ones in the training dataset. In this work, we use languages recognized by 3-state DFAs as these are the lowest nontrivial difficulty level, but this technique can be generalized to produce benchmarks of any difficulty level.
@@ 2.3 Language Understanding and LLMs
2.4 Training transformers on Formal Languages
Vafa et al. (2025) frame world modeling as a latent DFA identification task, finding that transformers trained on DFA traces (of massive DFAs representing board games and city maps) do not reconstruct the underlying DFA. Akyürek et al. (2024) find that transformers trained on 4-12 state DFA transducer traces more effectively learn to in-context-learn regular languages than RNNs. Here, we evaluate much larger and more capable LLMs on much simpler 3-state DFAs. However, our goal is evaluating the abilities of foundation models (trained on a broad corpus rather than a tailored one), and is thus not strictly easier.
@@ 3.1 DFAs and Regular Languages
The original Chomsky Hierarchy (Chomsky, 1959) separates language into four types (Figure 2). We focus on the task of understanding Type 3 languages, the simplest form of language in the hierarchy, that are recognized by a Deterministic Finite Automaton (DFA). Examples of languages recognized by DFAs include simple ones like binary strings with an even number of ones, and even such examples as numbers in base 10 divisible by 7. Type 3 languages are also known as regular languages, which are recognized by regular expressions.
One simple metric of the difficulty of a regular language is the number of states in the corresponding DFA, i.e., the amount of working memory. [^1] There are other metrics of difficulty, but we choose number of states as it is highly parsimonious. 2-state DFAs have the property that their set of states is no larger than the output set $\{0,1\}$, and, therefore, do not have any hidden state. We thus explore 3-state DFAs, as this is the simplest nontrivial case.
@@ 3.2 Sequence Completion Task
We first pose a sequence completion task, in which models must complete a sequence in a given DFA’s language. In practice, most language data that models encounter will be in roughly this format, with several example sequences in a given language followed by a distinct prefix that needs to be completed via next token prediction.
@@
We evaluate models by (1) sampling a DFA, (2) sampling 30 problem instances at random (each of which contains 30 example sequences and a distinct prefix), and then (3) computing a binary prediction score (whether or not the predicted completion creates a valid string in the language) for each instance separately, then computing a correctness metric as a fraction. We then average this metric over several sampled DFAs to produce our accuracy score.
@@ 3.3 Transducer Task
In this task, an input sequence is annotated with an output at each token, the final output is masked, and the masked output is predicted by a language model. We call this a transducer task, as the DFA converts a sequence of inputs into a sequence of outputs. E.g., given the language does the string have an even number of ’a’ tokens and the input abcabcaabbccaa, the annotated string (all that is provided to the model) is a0b0c0a1b1c1a0a1b1b1c1c1a0a and the output to predict is 1. For each problem instance, we provide 30 symbols, and for the first 29, the corresponding transducer output.
This task is significantly more transparent than the sequence completion task as the model has access to intermediate outputs, an (imperfect) proxy for intermediate state. n
@@ 4 Experiments
We produce versions of each of these prompts for each task, denoting these with a subscript S for sequence completion prompts and T for transducer prompts. Full listings of these prompts can be found in Appendix F. While no finite set of prompts will be fully sufficient to capture all possible model behavior, we believe our set of prompts captures common prompting strategies.
@@ 5.1 Sequence Completion
One potential problem with using this task for cross-model comparisons is the relevance of tokenization. Unfortunately, we found that forcing uniform tokenization by using commas in the prompt uniformly reduced accuracy, see Appendix E for details.
@@ 5.2 Transducer
Unlike sequence completion, this task is generally not fully determined, with the $\textsc{BruteForce}_{T}$ baseline only achieving 96.4% accuracy. Comparisons are still be valid as all models see the same fraction of unsolvable instances.
@@
We conclude that LLMs are unable to perform well on the DFA transducer inference task. This failure cannot be attributed to a lack of world modeling, as $n$-GramT models do not construct world models. Instead, it seems the LLMs are unable to detect patterns when those patterns are drawn from an unfamiliar source, even a relatively simple one.
@@ 5.4 Case Study: Sum Modulo 3 DFA
Total 100% 100%
a is no-op 70% 73%
1b and 1c lead to 0 47% 57%
2-periodic 30% 47%
3-periodic 13% 13%
2 red rooms 7% 10%
Figure 4: Results on Sum Modulo 3 DFA. (a) MB=mistral-nemo-minitron-8B/BasicT, CR=claude-3.5/Red-GreenT. Venn diagram of errors (out of 1000). Labeled percentages are accuracies. (b) Results of qualitative analysis, out of 30 in both cases.
@@
Figure 4a depicts the number of errors each model receives on 1000 instances of the transducer task for this DFA. Nearly all errors made by the 6-GramT model were also made by at least one LLM, while the two LLMs often made unique errors. While this task is better-known than most DFAs, all 3 models perform worse on this DFA than average.
We also performed a qualitative analysis, investigating CR’s outputs on the Red-GreenT prompt to see what kind of reasoning it is using; specifically we sampled 30 examples where it had the correct answer, and 30 examples where it had the incorrect answer but the 6-GramT model had the correct answer. Results of this analysis can be found in Figure 4b. We find that, in general, CR is following a 3-Gram approach, learning rules relating to the conditions under which the previous output and symbol can be used to predict the next output. Specifically, it is able to learn that a does not change the output, and that b and c will lead a 1 state to a 0 state. These results comport with the overall finding of Table 1, where we found that 3-GramT was the largest $n$-GramT that any LLM outperformed, as well as our finding that LLM performance decreases for tasks that are not solvable by $n$-Grams; see Appendix C for details.
@@ Appendix B Results of o1-preview
Appendix C Transducer results by difficulty class
Figure 5: Transducer results by difficulty class. We classify each DFA based on which of the baselines first achieves a score of 28/30 on the given instances. 6-Gram is excluded as it has very similar performance to 5-Gram. Each model’s best prompt results are plotted, with 95% confidence intervals, for all models with at least 100 DFAs; those with 10 or 30 had error bars too large to make this analysis useful.
Figure 5 displays results by difficulty level, as judged by the smallest $n$-Gram model that can solve a particular task. All models behave roughly monotonically, performing more poorly as difficulty increases. Additionally, we find that the best models continue to perform similarly to 4-Gram for tasks that 4-Gram does not perfectly solve.
Appendix D More details on Sum Modulo 3 DFA case study
@@
Appendix E Sequence Completion task prompt with Commas
@@
Appendix F Prompt Listings
@@ F.1 Summaries
F.2 Full example listings
F.2.1 BasicT